# Gravity versus astrophysics in black hole images and photon rings: equatorial emissions and spherically symmetric spacetimes

### v1

@@ 8.3 Conclusion
In this article, we have revisited the study already undertaken by Bauer et al. (2022) and Kocherlakota & Rezzolla (2022) on the prospect to test, via future interferometric observables of black holes like ngEHT (Johnson et al. 2023) or BHEX (Johnson et al. 2024), theories of gravity within the class of spherically symmetric RZ parametrised spacetimes (Rezzolla & Zhidenko 2014). Our main contributions to the effort is the inclusion of physically motivated synchrotron emission profiles in the equatorial plane (Vincent et al. 2022), parametrised according to the results of GRMHD simulations as in Desire et al. (2025), the simultaneous analysis of Keplerian, radial and mixed velocities of the thin accretion disc (Pu et al. 2016) and the simulations performed at the two higher observing frequencies of the ngEHT and BHEX, 230 and 345 GHz. We focus on the case of M87*, but our analysis could be generalised to other supermassive black-hole systems. Our findings can be summarised as follows:
@@
  • all the measured features of the direct image are too polluted by the astrophysical uncertainty on the complex accretion process to be used as a robust probe of the spacetime geometry, consistently with the results obtained for spherical accretion models (Bauer et al. 2022; Kocherlakota & Rezzolla 2022), so that present EHT observations, even if performed at higher observing frequencies, do not have the necessary resolution to detect a violation of general relativity without additional constraints on the properties of the accretion disc. Stated differently, the discriminatory test proposed in this paper requires access to at least the $n=1$ photon ring;
@@
  • the width of the $n=1$ intensity profiles along the polar angles on the screen are highly degenerate because the intrinsic emission profile of the disc governs their fall-off (Gralla et al. 2019); in other words, spacetime signatures on the width of the observed profiles can be mimicked by variations in the astrophysical configuration, so that the width of the $n=1$ photon ring is not a reliable quantity when looking for geometric modifications;
@@
  • the peak positions of the $n=1$ intensity profiles along the polar angles on the screen are the only investigated features that allow to disentangle the deviations induced by the spacetime from those produced by a class of viable emission profiles [^28] Within GR, the $n=1$ photon-ring radii are mainly influenced by the Kerr BH spin, rather than by astrophysical diversity (Desire et al. 2025). Besides, our assertion holds for some, but not all, alternative spherically symmetric BHs in Eichhorn et al. (2023), $n=2$ being otherwise needed., in agreement to the conclusions of preliminary works (Kocherlakota & Rezzolla 2022). Also, the degeneracy is broken unambiguously, if the mass over distance is well constrained and for all possible motions of the disc, ranging from a Keplerian to a radially-infalling one, when the synchrotron emissivity is computed including the value of the synchrotron pitch angle $\theta_{\mathrm{B}}$, and this is true for both components of a poloidal magnetic field. Since the velocity of realistic accreted matter is expected to have a predominant Keplerian fraction, this result is promising, although challenging, for M87* spacetime’s constraints via future very-long-baseline interferometry (VLBI) missions;
@@
  • Constraints on the black hole mass-to-distance estimate that are significantly more stringent than the present EHT ones, passing from an uncertainty on $M/D$ of about 11% (The Event Horizon Telescope Collaboration 2019d) to at least 2%, are needed to perform the discriminatory test of GR proposed in this paper using solely the intensity peak position of the first, or second, photon ring for small geometric deviations. On the other hand, some extreme metric deviations are already accessible with the current precision on $M/D$ given a sufficient resolving power;
@@
  • Relative measurements, independent of $M/D$, such as image asymmetry or images size ratio, cannot be used directly to distinguish a spherically-symmetric conservative spacetime [^29] Relative measurements are sufficient for non-parametrised alternatives to the Schwarzschild metric (Wielgus 2021; Eichhorn et al. 2023).. Yet, they are complementary probes of the astrophysical set-up, which is at present the main impediment to the identification of small observational deviations from GR.
@@ 8.4 Limitations of this work and perspectives
A series of approximations are present in our modelling and the results presented only stand under such simplifying hypotheses.
First of all, our equatorial disc is taken to be optically and geometrically thin, so that absorption effects along the geodesic motion are neglected, and our simulations do not include any turbulence or observing noise. However, the luminosity of M87*, about $3.6\times 10^{-6}L_{\mathrm{Edd}}$ (Prieto et al. 2016), where $L_{\mathrm{Edd}}$ is the Eddington luminosity, is very low, corresponding to a radiatively inefficient accretion flow (Narayan et al. 1995), and RIAFs nuclei are expected to be surrounded by a hot, geometrically thick disc (Yuan & Narayan 2014). The integration of the above mentioned elements could on one hand weaken and spoil the lensed features of the image and on the other hand add astrophysical uncertainty and, thus, degeneracy. Nevertheless, EHT data favour the scenario in which the emission is dominated by that of the equatorial plane and at 345 GHz absorption effects are reasonably weak (The Event Horizon Telescope Collaboration 2019c), so that the framework of this paper remains adapted to model future VLBI observations. In any case, it is the objective of future work to expand the query of this paper to more realistic geometrically thick discs, computing the full radiative transfer (Vincent et al. 2022) and incorporating statistical fluctuations (Lee & Gammie 2021).
Secondly, we have only varied here one parameter at a time, fixing the others at their Schwarzschild value. Nonetheless, particular parametric combinations, taken of course into their theoretically-allowed ranges (Kocherlakota & Rezzolla 2022), may reduce the deviations induced by the metric, increasing then the entanglement with the astrophysical possibilities. If it would be worth analysing the outputs of such choices, it would be even more compelling to create a new framework in which specific parametric deviations have clear physical interpretations.
Third, the discussions made here concern image-plane time-averaged images, but it would be of great interest to know if the obtained results could be translated in terms of interferometric signals and be exploited for a direct real-data analysis, see for example Cárdenas-Avendaño & Lupsasca (2023).
Finally, we have only considered spherically symmetric spacetimes, but realistic black holes are expected to be rotating. Since the spin of M87* is not well constrained, this would represent an additional astrophysical parameter to scan and which may mimic the observational divergences produced by metric deviations. Nevertheless, the relative measurements presented in Section 8.2 could be used to measure its value (Johnson et al. 2020; Broderick et al. 2022). A generalisation of the RZ metric to circular axisymmetric spacetimes already exists (Konoplya et al. 2016) and its implementation in a ray-tracing code could be straightforward.
@@ Appendix A Inclination effects on projections
To understand lensing band deformations caused by the inclination between an equatorial plane and a distant observer, it is useful to employ the bending number $w$ introduced in Section 2.
Let us consider the set-up of Figure A.1. The far-away observer’s yellow screen, labelled by the polar coordinates $(b,\psi)$ as in Figure 1, is situated at $\varphi=\varphi_{\infty}=0\bmod 2\pi$ rad. We define a $\psi$-direction on the observer’s screen as a line passing through its centre with a slope defined by the polar angle $\psi$. From the same arguments exposed in Section 2, it is immediate to see that photons reaching the same $\psi$-direction on the screen, belong to the same orbital plane. Let us also consider an infinitely large green plane cutting the black hole in two hemispheres and whose approaching side to the observer is directed towards the south of his screen [^30] Different orientations of the plane can be easily deduced by a corresponding rotation of the north of the screen.. The approaching side of the plane, rotated around the east-west axis, is inclined at an angle $i<\pi/2$ rad [^31] If the equatorial plane contained a disc, the latter is not seen edge-on. from the face-on grey plane. We call $\Sigma$ the red angle between the intersections, with the orbital plane, of the inclined plane and of the face-on plane. For the east-west $\psi=0$ direction on the screen, $\Sigma$ is zero because the pink plane passes precisely along the intersection of the green and grey ones. For the north-south $\psi=\frac{\pi}{2}$ direction, $\Sigma$ is equal to $i$, the relative inclination between the green and grey planes. When rotating the orbital plane from east to north, $\Sigma$ takes intermediate values; the orbital plane of Figure A.1 corresponds indeed to an intermediate $\psi$-direction, for which $0<\psi<\pi$.
Figure 10: Intersections between a face-on (grey) or inclined (green) plane, the yellow screen of a distant observer and a pink orbital plane.
@@
Figure 12: Null geodesics in a Schwarzschild spacetime reaching an observer on the far right. According to their winding number, geodesics can make up the inner shadow (brown), the $n=0$ image (orange) or the $n=1$ photon ring (light blue). We represent the projections of the black hole (black), the photon sphere (blue) and the equatorial plane for different polar directions on the screen: $\psi=0^{\circ}$ (pink), $45^{\circ}$ (red) and $90^{\circ}$ (violet).
Let us define the total change $\Delta\varphi$ of the orbital plane angle $\varphi$ and the bending number $w$ as in Gralla et al. (2019):
@@
Calling $\varsigma_{\mathrm{\lambda}}$ the sign of $\lambda$, we remark that $w=0$ for radial trajectories, $w\to\varsigma_{\mathrm{\lambda}}1/2$ when $b\to\infty$, while $w\to\varsigma_{\mathrm{\lambda}}\infty$ for $b\to\tilde{b}$. Having defined the bending number $w$ in equation (A2), we propose hereinbelow a classification of lensing bends which generalises previous conventions (Gralla et al. 2019; Chael et al. 2021) to arbitrary plane inclinations and polar directions on the observer’s screen. With $\varsigma_{\mathrm{\lambda}}$ the sign of $\lambda$ and $\sigma\coloneqq\Sigma/(2\pi)$, light rays are classified according to their order $n$, defined as follows: Thus, for $n\geq 1$, the $n$-th lensing band is the area on the plane of sky including all light rays whose bending number $w$ is inside the range settled by the order $n$, crossing then the equatorial plane exactly $n+1$ times along their way from the black hole to the observer or vice-versa, see Figure A.3. Given this definition, the periodicity of $\Sigma$, and so of $\sigma$, leads to the reflection symmetry of the lensing bands with respect to the north-south vertical direction on the observer’s screen. Put differently, if the projections of the equatorial plane in Figure A.3 tilt to the left from $\psi=0^{\circ}$ to $\psi=90^{\circ}$, they symmetrically come back to the initial perpendicular direction going towards $\psi=180^{\circ}$, imposing then the same geodesic orders.
@@ B.1 Asymptotes
Let us discuss the limits of $g$ at infinity and at the event horizon. At infinite distance from the black hole, it is legitimate to consider that the emitter is static and that the spacetime is flat, in other words $\boldsymbol{u}^{\mathrm{em}}=\boldsymbol{u}^{\mathrm{obs}}$ and $\boldsymbol{p}^{\mathrm{em}}=\boldsymbol{p}^{\mathrm{obs}}$. By means of equation (4.2), the redshift factor at radial infinity is:
@@
If the emitter is asymptotically close to the radius $r_{\mathrm{H}}$ of the event horizon, the radial component of its 4-velocity, $u^{r}_{\mathrm{em}}$, cannot be zero because below the radius of the ISCO, $r_{\mathrm{ms}}>r_{\mathrm{H}}$, no stable circular orbit is allowed and the 4-momentum of the emitted photon has to possess a radial component $p^{r}_{\mathrm{em}}$ to escape the gravitational attraction of the black hole and reach the observer at infinity. Besides, the 4-velocity of the emitter is proportional to its 4-momentum and 4-momenta’s time and azimuthal components are related to the constants of motion via equations (2.1). Then, at the horizon, the non-vanishing metric components of the RZ metric (3.2), (5), (6), with $x=1-r_{\mathrm{H}}/r$, read:
@@
Thus, the scalar product $\boldsymbol{p}^{\mathrm{em}}\cdot\boldsymbol{u}^{\mathrm{em}}=p^{\mathrm{em}}_{\,\mu}\,u_{\mathrm{em}}^{\,\mu}$, appearing at the denominator of equation (4.2), diverges because the constants of motion are finite and one can show that so are $p_{r}^{\mathrm{em}}$ and $u^{r}_{\mathrm{em}}$, by considering, respectively, the normalisation $p^{r}_{\mathrm{em}}p_{r}^{\mathrm{em}}=0$ and the photon’s geodesic equation along $\theta$ or the expression (4.1.2) of a fully radial infall. Thus, when $r$ tends to the radial position of the event horizon, for every equatorial velocity profile of the emitter $\boldsymbol{u}^{\mathrm{em}}$, the limit of the redshift factor is:
@@ B.2 Minkowski approximation
The Minkowski metric line element in spherical coordinates is:
@@
We recall that the redshift factor $g$ is given by (4.2):
@@
In Minkowski spacetime, the 4-momentum of the photon has a simple, geodesic invariant, analytical form (Vincent et al. 2024):
@@
where $\iota=\pi-i\in[0^{\circ};90^{\circ}]$ for $i\in[90^{\circ};180^{\circ}]$ and where the natural basis vectors $\boldsymbol{\partial}_{\mu}$ have been replaced by:
@@
As the velocity of a static observer is directly given by $\boldsymbol{u}^{\mathrm{obs}}=\mathbf{e_{t}}$, one has $\boldsymbol{p}^{\mathrm{obs}}\cdot\boldsymbol{u}^{\mathrm{obs}}=-1$ at the numerator of equation (B5). As for the scalar product appearing at the denominator of $g$ in equation (B5), it varies according to the 4-velocity of the emitter. The general expressions of a Keplerian and a radial 4-velocities are:
$&\boldsymbol{u}_{\text{Keplerian}}=u^{t}_{\text{Keplerian}}(\boldsymbol{\partial}_{t}+\Omega\boldsymbol{\partial}_{\varphi})\>,\quad&&\Omega=u^{\varphi}/u^{t}=d\varphi/dt\\ &\boldsymbol{u}_{\text{Radial}}=u^{t}_{\text{Radial}}(\boldsymbol{\partial}_{t}+\chi\boldsymbol{\partial}_{r})\>,\quad\ &&\>\chi=u^{r}/u^{t}=dr/dt\>.\end{aligned}$ (B8)
The expressions of $\Omega$ and $\chi$ depend on the spacetime metric and $u^{t}$ can be deduced by the normalisation $\boldsymbol{u}\cdot\boldsymbol{u}=-1$. If one takes the expressions of $\Omega$ and $\chi$ for the Schwarzschild spacetime, which coincide with the Newtonian results, one gets:
$\Omega&=\frac{1}{r}\sqrt{\frac{M}{r}}&&\Rightarrow\quad u^{t}_{\text{Keplerian}}=\sqrt{\frac{r}{r-M}}\\ \chi&=-\sqrt{\frac{2M}{r}}&&\Rightarrow\quad u^{t}_{\text{Radial}}=\sqrt{\frac{r}{r-2M}}\>.\end{aligned}$ (B9)
Thus, in the orthonormal basis $\mathbf{e}_{\mu}$, the 4-velocities read:
$&\boldsymbol{u}_{\text{Keplerian}}=\sqrt{\frac{r}{r-M}}\left(\mathbf{e}_{t}+\sqrt{\frac{M}{r}}\mathbf{e}_{\varphi}\right)\\ &\boldsymbol{u}_{\text{Radial}}=\sqrt{\frac{r}{r-2M}}\left(\mathbf{e}_{t}-\sqrt{\frac{2M}{r}}\mathbf{e}_{r}\right)\>.\end{aligned}$ (B10)
With the previous derivations, it is then immediate to get the expression of the denominator of $g$
$&(\boldsymbol{p}^{\mathrm{em}}\cdot\boldsymbol{u}^{\mathrm{em}})_{\text{Keplerian}}=-\sqrt{\frac{r}{r-M}}\left(1+\sqrt{\frac{M}{r}}\sin\iota\cos\varphi\right)\\ &(\boldsymbol{p}^{\mathrm{em}}\cdot\boldsymbol{u}^{\mathrm{em}})_{\text{Radial}}=-\sqrt{\frac{r}{r-2M}}\left(1-\sqrt{\frac{2M}{r}}\sin\iota\sin\varphi\right)\end{aligned}$ (B11)
and so the analytical formula of the redshift:
$g^{\text{Minkowski}}_{\text{Keplerian}}&=\left[\sqrt{\frac{r}{r-M}}\left(1+\sqrt{\frac{M}{r}}\sin\iota\cos\varphi\right)\right]^{-1}\\ g^{\text{Minkowski}}_{\text{Radial}}&=\left[\sqrt{\frac{r}{r-2M}}\left(1-\sqrt{\frac{2M}{r}}\sin\iota\sin\varphi\right)\right]^{-1}\>.\end{aligned}$ (B12)

### v2

@@ 8.3 Conclusion
In this article, we have revisited the study already undertaken by Bauer et al. (2022) and Kocherlakota & Rezzolla (2022) on the prospect to test, via future interferometric observables of black holes such as the ngEHT (Johnson et al. 2023) or BHEX (Johnson et al. 2024), theories of gravity within the class of spherically symmetric RZ parametrised space-times (Rezzolla & Zhidenko 2014). Our main contributions to the effort are the inclusion of physically motivated synchrotron emission profiles in the equatorial plane (Vincent et al. 2022), parametrised according to the results of GRMHD simulations as in Desire et al. (2025); the simultaneous analysis of Keplerian, radial, and mixed velocities of the thin accretion disc (Pu et al. 2016); and the simulations performed at the two higher observing frequencies of the ngEHT and BHEX, 230 and 345 GHz. Our focus has been on the case of M87*, but our analysis can be generalised to other supermassive black hole systems. Our findings are summarised as follows:
@@
  • All the measured features of the direct image are too polluted by the astrophysical uncertainty on the complex accretion process to be used as a robust probe of the space-time geometry. This is consistent with the results obtained for spherical accretion models (Bauer et al. 2022; Kocherlakota & Rezzolla 2022). Therefore, the present EHT observations, even if performed at higher observing frequencies, do not have the necessary resolution to detect a violation of general relativity without additional constraints on the properties of the accretion disc. Stated differently, the discriminatory test proposed in this paper requires access to at least the $n=1$ photon ring.
@@
  • The width of the $n=1$ intensity profiles along the polar angles on the screen are highly degenerate because the intrinsic emission profile of the disc governs their fall-off (Gralla et al. 2019). In other words, space-time signatures on the width of the observed profiles can be mimicked by variations in the astrophysical configuration, so the width of the $n=1$ photon ring is not a reliable quantity when looking for geometric modifications.
@@
  • The peak positions of the $n=1$ intensity profiles along the polar angles on the screen are the only investigated features that allow the deviations induced by the space-time to be disentangled from those produced by a class of viable emission profiles, [^28] Within general relativity, the $n=1$ photon-ring radii are mainly influenced by the Kerr black hole spin, rather than by astrophysical diversity (Desire et al. 2025). Besides, our assertion holds for some, but not all, alternative spherically symmetric black holes in Eichhorn et al. (2023), $n=2$ being otherwise needed. in agreement with the conclusions of preliminary works (Kocherlakota & Rezzolla 2022). Also, the degeneracy is broken unambiguously if the mass over distance is well constrained and for all possible motions of the disc, ranging from a Keplerian to a radially infalling velocity, when the synchrotron emissivity is computed including the value of the synchrotron pitch angle, $\theta_{\mathrm{B}}$. This is true for both components of a poloidal magnetic field. Since the velocity of realistic accreted matter is expected to have a predominant Keplerian fraction, this result is promising, although challenging, for M87* space-time’s constraints via future very-long-baseline interferometry (VLBI) missions.
@@
  • Constraints on the black hole mass-to-distance estimate that are significantly more stringent than the present EHT ones, passing from an uncertainty on $M/D$ of about 11% (The Event Horizon Telescope Collaboration 2019d) to at least 2%, are needed to perform the discriminatory test of general relativity proposed in this paper using solely the intensity peak position of the first, or second, photon ring for small geometric deviations. On the other hand, some extreme metric deviations are already accessible with the current precision on $M/D$ given a sufficient resolving power.
@@
  • Relative measurements, independent of $M/D$, such as image asymmetry or images size ratio, cannot be used directly to distinguish a spherically symmetric conservative space-time. [^29] Relative measurements are sufficient for non-parametrised alternatives to the Schwarzschild metric (Wielgus 2021; Eichhorn et al. 2023). Yet, they are complementary probes of the astrophysical set-up, which is at present the main impediment to the identification of small observational deviations from general relativity.
@@ 8.4 Limitations of this work and perspectives
A series of approximations are present in our modelling, and the results presented only stand under such simplifying hypotheses.
First of all, our equatorial disc was taken to be optically and geometrically thin so that absorption effects along the geodesic motion are neglected, and our simulations did not include any turbulence or observing noise. However, the luminosity of M87*, about $3.6\times 10^{-6}L_{\mathrm{Edd}}$ (Prieto et al. 2016), where $L_{\mathrm{Edd}}$ is the Eddington luminosity, is very low, corresponding to a radiatively inefficient accretion flow (Narayan et al. 1995), and RIAFs nuclei are expected to be surrounded by a hot geometrically thick disc (Yuan & Narayan 2014). The integration of the above mentioned elements could, on one hand, weaken and spoil the lensed features of the image and, on the other hand, add astrophysical uncertainty and thus degeneracy. Nevertheless, EHT data favour the scenario in which the emission is dominated by that of the equatorial plane, and at 345 GHz absorption effects are reasonably weak (The Event Horizon Telescope Collaboration 2019c), so the framework of this paper remains adapted to model future VLBI observations. In any case, it is the objective of future work to expand the query of this paper to more realistic geometrically thick discs, compute the full radiative transfer (Vincent et al. 2022), and incorporate statistical fluctuations (Lee & Gammie 2021).
Secondly, here we only varied one parameter at a time, fixing the others at their Schwarzschild value. Nonetheless, particular parametric combinations, taken of course into their theoretically allowed ranges (Kocherlakota & Rezzolla 2022), may reduce the deviations induced by the metric, then increasing the entanglement with the astrophysical possibilities. If it would be worth analysing the outputs of such choices, it would be even more compelling to create a new framework in which specific parametric deviations have clear physical interpretations.
Third, the discussions made here concern image-plane time-averaged images, but it would be of great interest to know if the obtained results could be translated in terms of interferometric signals and be exploited for a direct real-data analysis (see for example Cárdenas-Avendaño & Lupsasca (2023)).
Finally, we only considered spherically symmetric space-times, but realistic black holes are expected to be rotating. Since the spin of M87* is not well constrained, this would represent an additional astrophysical parameter to scan, and the spin may mimic the observational divergences produced by metric deviations. Nevertheless, the relative measurements presented in Section 8.2 could be used to measure its value (Johnson et al. 2020; Broderick et al. 2022). A generalisation of the RZ metric to circular axisymmetric space-times already exists (Konoplya et al. 2016), and its implementation in a ray-tracing code could be straightforward.
@@ Appendix A Inclination effects on projections
To understand lensing band deformations caused by the inclination between an equatorial plane and a distant observer, it is useful to employ the bending number, $w$.
We considered the set-up of Figure A.1. The far-away observer’s yellow screen, labelled by polar coordinates, $(b,\psi)$, as in Figure 1, is situated at $\varphi=\varphi_{\infty}=0\bmod 2\pi$ rad. We defined a $\psi$-direction on the observer’s screen as a line passing through its centre with a slope defined by the polar angle, $\psi$. From the same arguments exposed in Section 2, it is immediate to see that photons reaching the same $\psi$-direction on the screen, belong to the same orbital plane. We also considered an infinitely large green plane cutting the black hole into two hemispheres and whose approaching side to the observer is directed towards the south of their screen. [^30] Different orientations of the plane can be easily deduced by a corresponding rotation of the north of the screen. The approaching side of the plane, rotated around the east-west axis, is inclined at an angle of $i<\pi/2$ rad [^31] If the equatorial plane contained a disc, the latter is not seen edge-on. from the face-on grey plane. We called $\Sigma$ the red angle between the intersections, with the orbital plane, of the inclined plane and of the face-on plane. For the east-west $\psi=0$ direction on the screen, $\Sigma$ is zero because the pink plane passes precisely along the intersection of the green and grey ones. For the north-south $\psi=\frac{\pi}{2}$ direction, $\Sigma$ is equal to $i$, the relative inclination between the green and grey planes. When rotating the orbital plane from east to north, $\Sigma$ takes intermediate values; the orbital plane of Figure A.1 corresponds indeed to an intermediate $\psi$-direction, for which $0<\psi<\pi$.
Figure 10: Intersections between a face-on (grey) or inclined (green) plane, the yellow screen of a distant observer, and a pink orbital plane.
@@
Figure 12: Null geodesics in a Schwarzschild space-time reaching an observer on the far right. According to their winding number, geodesics can make up the inner shadow (brown), the $n=0$ image (orange), or the $n=1$ photon ring (light blue). We represent the projections of the black hole (black), the photon sphere (blue), and the equatorial plane for different polar directions on the screen: $\psi=0^{\circ}$ (pink), $45^{\circ}$ (red), and $90^{\circ}$ (violet).
We defined the total change, $\Delta\varphi$, of the orbital plane angle, $\varphi$, and the bending number, $w$, as in Gralla et al. (2019):
@@
Calling $\varsigma_{\mathrm{\lambda}}$ the sign of $\lambda$, we remark that $w=0$ for radial trajectories, $w\to\varsigma_{\mathrm{\lambda}}1/2$ when $b\to\infty$, while $w\to\varsigma_{\mathrm{\lambda}}\infty$ for $b\to\tilde{b}$. Having defined the bending number $w$ in equation (A2), we propose hereinbelow a classification of lensing bands that generalises previous conventions (Gralla et al. 2019; Chael et al. 2021) to arbitrary plane inclinations and polar directions on the observer’s screen. With $\varsigma_{\mathrm{\lambda}}$ the sign of $\lambda$ and $\sigma\coloneqq\Sigma/(2\pi)$, light rays are classified according to their order, $n$, defined as follows: Thus, for $n\geq 1$, the $n$-th lensing band is the area on the plane of sky including all light rays whose bending number, $w$, is inside the range settled by the order $n$, crossing then the equatorial plane exactly $n+1$ times along their way from the black hole to the observer or vice-versa (see Figure A.3). Given this definition, the periodicity of $\Sigma$, and so of $\sigma$, leads to the reflection symmetry of the lensing bands with respect to the north-south vertical direction on the observer’s screen. Put differently, if the projections of the equatorial plane in Figure A.3 tilt to the left from $\psi=0^{\circ}$ to $\psi=90^{\circ}$, they symmetrically come back to the initial perpendicular direction going towards $\psi=180^{\circ}$, imposing then the same geodesic orders.
@@ B.1 Asymptotes
Here, we discuss the limits of $g$ at infinity and at the horizon. At infinite distance from the black hole, it is legitimate to consider that the emitter is static and that the space-time is flat, in other words $\boldsymbol{u}^{\mathrm{em}}=\boldsymbol{u}^{\mathrm{obs}}$ and $\boldsymbol{p}^{\mathrm{em}}=\boldsymbol{p}^{\mathrm{obs}}$. By means of equation (4.2), the redshift factor at radial infinity is
@@
If the emitter is asymptotically close to the radius of the event horizon, $r_{\mathrm{H}}$, the radial component of its 4-velocity, $u^{r}_{\mathrm{em}}$, cannot be zero because below the radius of the ISCO, $r_{\mathrm{ms}}>r_{\mathrm{H}}$, no stable circular orbit is allowed, and the 4-momentum of the emitted photon has to possess a radial component, $p^{r}_{\mathrm{em}}$, to escape the gravitational attraction of the black hole and reach the observer at infinity. Besides, the 4-velocity of the emitter is proportional to its 4-momentum and 4-momenta’s time and azimuthal components are related to the constants of motion via equations (2.1). Then, at the horizon, the non-vanishing metric components of the RZ metric (3.2), (5), (6), with $x=1-r_{\mathrm{H}}/r$, read
@@
Thus, the scalar product $\boldsymbol{p}^{\mathrm{em}}\cdot\boldsymbol{u}^{\mathrm{em}}=p^{\mathrm{em}}_{\,\mu}\,u_{\mathrm{em}}^{\,\mu}$, appearing at the denominator of equation (4.2), diverges because the constants of motion are finite and one can show that so are $p_{r}^{\mathrm{em}}$ and $u^{r}_{\mathrm{em}}$, by considering, respectively, the normalisation $p^{r}_{\mathrm{em}}p_{r}^{\mathrm{em}}=0$ and the photon’s geodesic equation along $\theta$ or the expression (4.1.2) of a fully radial infall. Thus, when $r$ tends to the radial position of the event horizon, for every equatorial velocity profile of the emitter, $\boldsymbol{u}^{\mathrm{em}}$, the limit of the redshift factor is
@@ B.2 Minkowski approximation
The Minkowski metric line element in spherical coordinates is
@@
We recall that the redshift factor, $g$, is given by (4.2)
@@
In Minkowski space-time, the 4-momentum of the photon has a simple, geodesic invariant, analytical form (Vincent et al. 2024):
@@
where $\iota=\pi-i\in[0^{\circ};90^{\circ}]$ for $i\in[90^{\circ};180^{\circ}]$ and where the natural basis vectors, $\boldsymbol{\partial}_{\mu}$, have been replaced by
@@
As the velocity of a static observer is directly given by $\boldsymbol{u}^{\mathrm{obs}}=\mathbf{e_{t}}$, one has $\boldsymbol{p}^{\mathrm{obs}}\cdot\boldsymbol{u}^{\mathrm{obs}}=-1$ at the numerator of equation (B5). As for the scalar product appearing at the denominator of $g$ in equation (B5), it varies according to the 4-velocity of the emitter. The general expressions of a Keplerian and a radial 4-velocities are
$\begin{aligned} &\boldsymbol{u}_{\text{Keplerian}}=u^{t}_{\text{Keplerian}}(\boldsymbol{\partial}_{t}+\Omega\boldsymbol{\partial}_{\varphi})\>,\quad&&\Omega=u^{\varphi}/u^{t}=d\varphi/dt\\ &\boldsymbol{u}_{\text{Radial}}=u^{t}_{\text{Radial}}(\boldsymbol{\partial}_{t}+\chi\boldsymbol{\partial}_{r})\>,\quad\ &&\>\chi=u^{r}/u^{t}=dr/dt\>.\end{aligned}$ (B8)
The expressions of $\Omega$ and $\chi$ depend on the space-time metric and $u^{t}$ can be deduced by the normalisation $\boldsymbol{u}\cdot\boldsymbol{u}=-1$. If one takes the expressions of $\Omega$ and $\chi$ for the Schwarzschild space-time, which coincide with the Newtonian results, one gets
$\begin{aligned} \Omega&=\frac{1}{r}\sqrt{\frac{M}{r}}&&\Rightarrow\quad u^{t}_{\text{Keplerian}}=\sqrt{\frac{r}{r-M}}\\ \chi&=-\sqrt{\frac{2M}{r}}&&\Rightarrow\quad u^{t}_{\text{Radial}}=\sqrt{\frac{r}{r-2M}}\>.\end{aligned}$ (B9)
Thus, in the orthonormal basis $\mathbf{e}_{\mu}$, the 4-velocities read
$\begin{aligned} &\boldsymbol{u}_{\text{Keplerian}}=\sqrt{\frac{r}{r-M}}\left(\mathbf{e}_{t}+\sqrt{\frac{M}{r}}\mathbf{e}_{\varphi}\right)\\ &\boldsymbol{u}_{\text{Radial}}=\sqrt{\frac{r}{r-2M}}\left(\mathbf{e}_{t}-\sqrt{\frac{2M}{r}}\mathbf{e}_{r}\right)\>.\end{aligned}$ (B10)
With the previous derivations, it is then immediate to obtain the expressions of the denominator of $g$:
$\begin{aligned} &(\boldsymbol{p}^{\mathrm{em}}\cdot\boldsymbol{u}^{\mathrm{em}})_{\text{Keplerian}}=-\sqrt{\frac{r}{r-M}}\left(1+\sqrt{\frac{M}{r}}\sin\iota\cos\varphi\right)\\ &(\boldsymbol{p}^{\mathrm{em}}\cdot\boldsymbol{u}^{\mathrm{em}})_{\text{Radial}}=-\sqrt{\frac{r}{r-2M}}\left(1-\sqrt{\frac{2M}{r}}\sin\iota\sin\varphi\right)\end{aligned}$ (B11)
and the analytical formulae of the redshift are thus
$\begin{aligned} g^{\text{Minkowski}}_{\text{Keplerian}}&=\left[\sqrt{\frac{r}{r-M}}\left(1+\sqrt{\frac{M}{r}}\sin\iota\cos\varphi\right)\right]^{-1}\\ g^{\text{Minkowski}}_{\text{Radial}}&=\left[\sqrt{\frac{r}{r-2M}}\left(1-\sqrt{\frac{2M}{r}}\sin\iota\sin\varphi\right)\right]^{-1}\>.\end{aligned}$ (B12)